OpenStudy (anonymous):

divide, answer a+bi form: (2+8i)/(7+2i)

OpenStudy (anonymous):

\[(\ \frac{2 + 8i}{7+2i})(\ \frac{7 - 2i}{7-2i}) = [14 - 4i + 56i - 16i ^{2}]/[49 - 4i ^{2}\]]

OpenStudy (anonymous):

As we perform arithmetic operations note that i^2 = -1

OpenStudy (anonymous):

Now we add like terms: \[(\ \frac{30 + 54i }{53})\]

OpenStudy (anonymous):

Since i^2 = -1 , -16i^2 = 16 16 + 14 = 30 Also 56i - 4i = 54i (*Because they are like terms we can add them)

OpenStudy (anonymous):

So the numerator simplifies to (30 + 52i) (*52i sorry because its (56 - 4)

OpenStudy (anonymous):

For the denominator we have 49 - 4i^2 = 49 + 4 = 53

OpenStudy (anonymous):

The correct final answer is therefore \[(\ \frac{30 + 52i}{53})\]

jimthompson5910 (jim_thompson5910):

Answer is \[\large \frac{2+8i}{7+2i}=\frac{30}{53}+\frac{52}{53}i\] See attached for work shown.

myininaya (myininaya):

jim did you type that yourself?